OpenStudy (anonymous):

f(x)=5x^4-(1/3x^2)+ax f'(x)=?

OpenStudy (ajprincess):

Differentiation of f(x) with respect to x gives f'(x) Is ur question \[\frac{5x^4-1}{3x^2+ax}\]?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

1/3(x)

OpenStudy (ajprincess):

\(f(x)=bx^n\) \(f'(x)=b*nx^{n-1}\) does this help?

OpenStudy (anonymous):

i would have to apply on all those terms

OpenStudy (ajprincess):

f(x)=b*x^n. this is similar to f(x)=5x^4 in place of b u have 5 and n is 4. f'(x) of bx^n is b*nx^{n-1}. so can u find f'(x) of 5x^4.

OpenStudy (ajprincess):

u can use it for other terms (-1/3)x^2 and +ax. Can u find @jadi?

OpenStudy (anonymous):

i would have to apply more then one time ???

OpenStudy (ajprincess):

yup. hav u found the differentiation of 5x^4?

OpenStudy (anonymous):

20x^3

OpenStudy (ajprincess):

ya. very good:) similarly find differentiation of (-1/3)x^2 and +ax and combine them to find the total differentiation of f(x)=5x^4-(1/3x^2)+ax

OpenStudy (ajprincess):

@jadi hav u found?

OpenStudy (anonymous):

it 0

OpenStudy (ajprincess):

noo. what is the differentiation of ax?

OpenStudy (anonymous):

a

OpenStudy (ajprincess):

ya good:) and diffrentiation of (-1/3)x^2?

OpenStudy (anonymous):

2/3x

OpenStudy (ajprincess):

Guess u forgot the - sign. Now add them all together to find the differentiation of 5x^4-(1/3x^2)+ax. Can u do @jadi?

OpenStudy (anonymous):

no i could not

OpenStudy (ajprincess):

f(x)=5x^4-(1/3x^2)+ax f'(x)=20x^3-(2/3)x+a Getting it?

OpenStudy (anonymous):

ja is that all

OpenStudy (ajprincess):

yup:)

OpenStudy (anonymous):

ooo thank u very much

OpenStudy (ajprincess):

welcome:)